A sample of size n=10 is drawn from the population. The data is shown below.

22. A sample of size n = 10 is drawn from a population. The data is shown below. 87.3 71.4 67.6 58.6 60.6 55.4 87.3 49.2 87.3 66.2 What is the range of this data set? range What is the standard deviation of this data set? (Remember, it is a sample.) Please report the answer with appropriate rounding, reporting 2 more decimal places than the original data. Please. please please do not calculate the value by hand stdev
Answer #1

a) Range= Max- Min= 87.3-49.2 =38.1

b)

X (X-X_bar)^2
67.6 2.2201
87.3 331.6041
58.6 110.0401
60.6 72.0801
55.4 187.4161
87.3 331.6041
71.4 5.3361
49.2 395.6121
87.3 331.6041
66.2 8.3521
Mean 69.09
Sum 1775.869

Mean: \sum X_i/N =69.09

Standard Deviation: s= \sqrt{\sum (X_i-Mean)^2/N-1} =\sqrt{\frac{1775.869}{9}}= 14.04702

stdev= 14.04702